$SPAD/src/input ribbon.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
--the first attempt
drawRibbons(flist,xrange) ==
sp := createThreeSpace()
y0 := 0
for f in flist repeat
  makeObject(f,xrange,y=y0..y0+1,
     space==sp, var2Steps ==1)
  y0 := y0+1
vp:=makeViewport3D(sp,"Ribbons")
drawStyle(vp,"shade")
outlineRender(vp,"on")
showRegion(vp,"on")
n := # flist
zoom(vp,n,1,n)
rotate(vp,0,75)
vp

———-

References^{[1] nothing}

4